Elasticity of demand for Good X is half of elasticity of demand for Good Y. Quantity demand for
Good Y falls by 15 percent when its price rises by Rs.5 per unit from Rs. 10 per unit. Calculate
percentage change in demand for Good X if its price falls by 50 percent.
1
Expert's answer
2020-09-21T08:19:29-0400

"Edy = \frac{\Delta Dy} {\Delta Py}
 = \frac{15} {50}
 = 0.3"


"Edx = 0.5(Edy)\n = 0.5 (0.3)\n = 0.15"


"Edx = \frac{\Delta Dx} {\Delta Px}"


"0.15 = \frac{\Delta Dx} {50}"


"\Delta Dx = 0.15(50) = 7.5"


Therefore, the demand for Good X changes by 7.5 percent if its price falls by 50 percent